OpenStudy (calculusxy):

Please Check My Answer! The graph of y = 2/3x - 12 intersects the y-axis at Q(0,-12) and is perpendicular to a line joining Q to the point P(x, 0).  Find x. My answer: x = -8

OpenStudy (calculusxy):

@Nnesha

OpenStudy (calculusxy):

Am i correct or not?

Nnesha (nnesha):

any work??? :=))

OpenStudy (calculusxy):

Yes.

imqwerty (imqwerty):

try finding out the slope of line joining (-8,0) and (0,-12) then check if its appropriate or not :) u'll find out if ur correct or not

OpenStudy (calculusxy):

Perpendicular Slopes = opposite reciprocal \[\large y = mx + b \] y = y-coordinate m = slope x = x-coordinate b = y-intercept Our equation: \[\large y = \frac{ 2 }{ 3 }x - 12\] Slope: \(\large \frac{2}{3}\) Perpendicular Slope is \(\large -\frac{3}{2}\) New equation with perpendicular slope: \[\large y = -\frac{ 3 }{ 2 }x - 12\] We have the coordinate: \(\large (x, 0)\) where \(0\) is the y-coordinate and \(x\) is the unknown x-coordinate. \(\large 0 = -\frac{3}{2} x- 12\) Solve for x to get: \(\large -8\)

Nnesha (nnesha):

that's correct. to double check you can find slope of two points as mentioned by imqwerty

OpenStudy (calculusxy):

okay thanks :) i have another question can you please check my answer: A line with x-intercept -4 passes through (2,6) and (p,10).  Find p. answer: p = 6

Nnesha (nnesha):

check your answer let p=6 (2,6)(6,10) find slope and y-intercept and then substitute y for zero to find x-intercept if you get -4 as final answer then yes that is correct value for p